A sample of chalk (caco3) is contaminated with calcium sulphate (caso4); 1 gm of the solid mixture is dissolved in 230 ml of N/10 HCl; 40 ml of N/10 NaOH is required to neutralise the excess of acid. Find the % of chalk in the mixture.

Grade:11

1 Answers

Yashee Sinha
6 years ago
The truck to this question is the fact that NaOH is required to neutralize *excess* of acid.That implies that,Gram equivalents of the base (NaOH) = Gram equivalents of the excess of acid (HCl)(40/1000) × (1/10) = excess of Gram equivalents of HCl = 0.004=> Gram equivalents that neutralized CaCO3 = Gram equivalents of HCl - 4 = (0.230 × 1/10) - 0.004 = 0.019=> Gram equivalents of CaCO3 = 0.019=> Mass of CaCO3 = 0.019 × 50 (equivalent mass = 50) = 0.95So, % of CaCO3 = 0.95/1 × 100 = 95%